# Sina Baharlouei

### 06 Dec 2023 v1

@@ 3.2 Robust $f$-FERM Under $\ell_{\infty}$ Norms and Potentially Large Distribution Shifts
For the general $f$-divergence in (8), it is easy to show that the function $D_{f}$ is convex in $\mathbb{P}$ and $\mathbb{Q}$. Thus, under $\ell_{\infty}$ uncertainty set on $\mathbb{P}$ and $\mathbb{Q}$, the optimal solution of the maximization problem in (8) will be at an extreme point. Moreover, under the assumption that $f^{*}(\alpha^{*})>0$ and $\alpha^{*}>0$ (which is satisfied for KL divergence), one can easily see that the optimal $p_{j}=\min\{\hat{p}_{j}+\delta,1\}$ and $q_{j}=\max\{\hat{q_{j}}-\delta,0\}$ (see Appendix I for the exact proof). Notice that to obtain this efficient optimal closed-form solution, we need to relax the probability simplex constraint. Thus under this assumption, problem (8) can be reformulated as
$\min_{\bm{\theta}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\mathcal{D}_{f}(\min\{\mathbb{P}+\delta,1\}||\max\{\mathbb{Q}-\delta,0\}),$ (12)
which is a regular minimization problem, and (sub)gradient descent can be utilized to solve it.
@@ 4 Experiments
This section evaluates the performance and efficiency of the proposed $f$-divergence frameworks in Section 2 and its robust version against distribution shifts (Section 3) through extensive experiments on benchmark datasets against several state-of-the-art approaches. We use $3$ popular notions of group fairness, including demographic parity, equalized odds, and equality of opportunity violations (see Appendix A for the exact definitions) to measure the fairness of trained models. To run Algorithm 1, we set $\eta_{{\bm{\theta}}}$ and $\eta_{\alpha}$ to $10^{-5}$ and $10^{-6}$ respectively in all experiments. Further, by changing $\lambda$, we get different points in the trade-off curve between the accuracy and fairness of the trained model. The range of values for $\lambda$ depends on the $f$-divergence we use (see Appendix J for more information on tuning hyper-parameters). In the inference phase of our experiments, we use the standard maximum likelihood decoding based on the output of the softmax layer, i.e., the predicted label is the label with the highest logit value.
@@ 4.1 Fairness-Accuracy Tradeoffs on Benchmark Datasets
Figure 1: Performance of different $f$-divergences as the regularizers for measuring fainress violation. The experiment is on the adult dataset with gender and race as sensitive attributes. While the offered tradeoffs are close to each other for small demographic parity violations, KL-divergence shows an extraordinary performance for a low-fairness high-accuracy regime. We do not display the performance for larger batch sizes or when only one sensitive attribute is available due to the insignificant difference between the performance of different $f$-divergences.
In the first set of experiments, we compare different $f$-divergence formulations for ($f$-FERM) to each other and several state-of-the-art approaches supporting multiple sensitive attributes. Figure 1 demonstrates the given tradeoff on the adult dataset (Becker and Kohavi, 1996) with gender and race as the sensitive attributes (black-female, black-male, white-female, white-male). To measure fairness, we use the demographic parity violation defined as:
$\textrm{DPV}=\max_{i,j\in\mathcal{S}}|\mathbb{P}(\hat{y}=1|s=i)-\mathbb{P}(\hat{y}=1|s=j)|$
In the case of binary sensitive attributes (e.g., gender), there is no significant variation between different $f$-divergences. However, when we have $2$ sensitive attributes and the batch size is small ($8$ in Figure 1), the results significantly differ for various $f$-divergences. Interestingly, KL-divergence for smaller $\lambda$ values shows improvement in fairness violation and accuracy simultaneously. We do not observe such a phenomenon for other $f$-divergences and state-of-the-art approaches in the literature. Further, in Figure 2, we compare one of the $f$-divergences (reverse KL) to several SOTA methods including Mary et al. (2019); Baharlouei et al. (2020); Cho et al. (2020). Other approaches such as the pre-processing method of Zemel et al. (2013), post-processing approach of Hardt et al. (2016), and several in-processing methods including Zafar et al. (2017); Donini et al. (2018); Jiang et al. (2020) demonstrate lower performance compared to the ones depicted in Figure 2 and are removed from the figure. While our approach demonstrates consistently good performance across different batch sizes (full-batch, $64$, $8$, $2$), other methods’ performances drop significantly for smaller ones. For further experiments on other datasets (German and COMPAS) and other fairness measures (equality of opportunity and equalized odds violations), see Appendix K.
@@ 4.2 Fairness-Accuracy Tradeoffs in the Presence of the Distribution Shift
We consider two experiments to evaluate the robustness of Algorithms developed in Section 3. In the first experiment, we randomly switch the label of genders for $n\%$ of the data points ($n$ ranges from $1$ to $20$) in the Adult dataset. Then, we train models on the new datasets with a proportion of corrupted sensitive attributes and evaluate the performance on the test data. Figure 3 is obtained by training different models to achieve $80\%$ accuracy on the test data and comparing their demographic parity violation. By increasing the percentage of corrupted sensitive attributes, we see that both $f$-DRO and $f$-infinity achieve less DP violation than SOTA approaches in the literature. In this specific experiment, $f$-DRO works better than $f$-infinity, and there is no significant difference between choosing KL-divergence or $\chi^{2}$ as the function $f$. Among the papers designed for handling distribution shifts, Rezaei et al. (2021) and Wang et al. (2020) were the only options with the available implementation.
@@
In a more recently collected dataset (new adult) (Ding et al., 2021), the users are separated based on their living state. We train different fair models in a single state and evaluate the fairness-accuracy tradeoff in other states. Figure 4 depicts the performance of different methods. For each method, the center point is the average of accuracy and fairness among $50$ states. The horizontal and vertical lines show the $25$-percentile to $75$-percentile range of performance among the states. The training fairness violation is set to $0.02$ for all methods. We observe that $f$-infinity preserves the fairness level better than other approaches. In comparison, $f$-DRO has a better accuracy. Depending on the application, we suggest using $f$-infinity if preserving a high level of fairness is a priority and $f$-DRO for the cases when a better tradeoff between fairness and accuracy is expected. Note that both of these approaches offer better fairness-accuracy tradeoffs compared to the SOTA approaches in the literature.
@@ Appendix A $f$-FERM for other notions of group fairness
This section shows how we can use alternative notions of fairness, such as equality of opportunity of equalized odds [Hardt et al., 2016] instead of demographic parity violation in $f$-FERM.
@@
Therefore, we must add a regularizer per each class label to satisfy the equalized odds notion. Other notions of fairness can be used in this framework as long as they can be represented as the conditional independence between sensitive attributes, predictions, and labels [Castelnovo et al., 2022].
@@ Appendix B $f$-divergences for continuous sensitive attributes and target variables
In Section 2, we developed a framework for promoting fairness for classification problems where both target labels and sensitive attributes are discrete variables. Hence we could efficiently solve the variational formulation that arose through the designing of unbiased estimators. However, it is not uncommon to find applications of f-divergence regularizers in practice that require either the sensitive features or the output variable to be continuous; or both to be continuous parameters. In such cases, the summation over the respective variable is replaced by an integral over the probability distribution. The challenging aspect is calculating the variational form’s integral and trailing supremum in the continuous domain.
@@ Appendix C $f$-divergences cover well-known notions of fairness violation
Proposition 4
[Polyanskiy and Wu, 2022, Theorem 2.3] Let $f$ be a convex function from $\mathbb{R}^{+}$ to $\mathbb{R}$, such that $f$ is convex, $f(1)=0$, $f$ is strictly convex in a neighborhood of $1$. Then $\mathcal{D}_{f}(\mathbb{P}||\mathbb{Q})=0$, if and only if $P=Q$.
@@
Proposition 5
Exponential Rényi Mutual Information (ERMI) [Mary et al., 2019, Lowy et al., 2022] is an $f$-divergence with $f(t)=(t-1)^{2}$
Proof  Exponential Rényi Mutual Information is defined as [Lowy et al., 2022]:
@@
Proof  Based on Propositions 5, ERMI is an $f$-divergence with $f(t)=(t-1)^{2}$. Therefore, the proposition is an immediate result of Lemma 3 in [Lowy et al., 2022].
Proposition 7
Rényi correlation [Baharlouei et al., 2020] can be upper bounded by the $f$-divergences for the choice of $f(t)=(t-1)^{2}$.
Proof  Based on Propositions 5, ERMI is an $f$-divergence with $f(t)=(t-1)^{2}$. Therefore, the proposition is an immediate result of Lemma 2 in [Lowy et al., 2022].
Remark 8
Mutual Information as the measure of fairness violation [Cho et al., 2020] is a special case of $f$-divergences for the choice of KL-divergence $f(t)=t\log(t)$ in ($f$-FERM).
Appendix D Proof of Proposition 1
Lemma 9
@@ Appendix D Proof of Proposition 1
Proof  Let $g$ be the convex conjugate of the function $f$ defined as:
@@
Since $f$ is a lower semi-continuous convex function, by Fenchel-Moreau theorem [Ioffe and Tihomirov, 2009], it is biconjugate, which means the taking conjugate of $g$ transforms it back to $f$. Therefore,
@@
where $g$ is the convex conjugate of $f$.
@@
Set $\mathbb{P}=\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s}),\mathbb{Q}=\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s),$ and $\alpha_{i}={\mathbf{A}}_{jk}$. Therefore, we obtain the formulation in (3).
@@ Appendix E Derivation of Closed-Form Expressions for Unbiased Gradient Estimators of $f$-Divergences
Proof  Proof follows naturally from [Polyanskiy and Wu, 2022, Proposition 7.2]
Theorem 11
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}A_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)-({\mathbf{A}}_{jk}+\frac{{\mathbf{A}}_{jk}^{2}}{4})\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (19)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}{\mathbf{A}}_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)+(1+\ln(-{\mathbf{A}}_{jk}))\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (20)
@@
Theorem 13
Let $f(t)=\frac{1}{2}|t-1|$ and $\mathbb{P}(s=k)=\pi_{k}$ (Total Variational Distance). Then, Equation (1) can be written as (where $|{\mathbf{A}}_{jk}|\leq\frac{1}{2}$):
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}{\mathbf{A}}_{jk}\Big{[}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)-\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (21)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}{\mathbf{A}}_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)-e^{{\mathbf{A}}_{jk}-1}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (22)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}{\mathbf{A}}_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)+\ln(2-e^{{\mathbf{A}}_{jk}})\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (23)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}{\mathbf{A}}_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)\\ -\frac{\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)}{\alpha}\Big{(}\Big{(}(\alpha-1){\mathbf{A}}_{jk}+1\Big{)}^{\frac{\alpha}{\alpha-1}}-1\Big{)}\Big{]}$ (24)
@@
This is tantamount to the substitution $\alpha\leftarrow\frac{1+\alpha}{2}$ in the original definition of generalized f-divergence.
Theorem 17
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}{\mathbf{A}}_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)+\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{(}\frac{1}{{\mathbf{A}}_{jk}}+2\Big{)}\Big{]}$ (25)
@@
Appendix F Formal Statement of Theorem 2 and Proof
Theorem 18
@@ Appendix F Formal Statement of Theorem 2 and Proof
Now, we are ready to prove Theorem 18.
Proof  The proof consists of three main steps. First, we must show that the gradient estimator in Algorithm 1 is unbiased. Since the samples are IID, for any function $\psi(\cdot,\cdot)$, and an IID batch of data points $\mathcal{B}$ we have:
@@
$\min_{{\bm{\theta}}}\>\max_{{\mathbf{A}}}\quad\frac{1}{n}\sum_{i=1}^{n}\left[\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}{\mathbf{A}}_{jk}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}({\mathbf{A}}_{jk})\pi_{k}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\Big{]}\right]$ (27)

### 07 Apr 2024 v2

@@ 3.2 Robust $f$-FERM Under $\ell_{\infty}$ Norms and Potentially Large Distribution Shifts
Since the function $D_{f}$ is convex in $\mathbb{P}$ and $\mathbb{Q}$, under $\ell_{\infty}$ uncertainty set on $\mathbb{P}$ and $\mathbb{Q}$, the optimal solution of the maximization problem in (8) will be at an extreme point. Moreover, under the assumption that $f^{*}(\alpha^{*})>0$ and $\alpha^{*}>0$ (which is satisfied for KL divergence), one can easily see that the optimal $p_{j}=\min\{\hat{p}_{j}+\delta,1\}$ and $q_{j}=\max\{\hat{q_{j}}-\delta,0\}$ (see Appendix I for the exact proof). Notice that we need to relax the probability simplex constraint to obtain this efficient, optimal closed-form solution. Thus under this assumption, problem (8) can be reformulated as
$\vspace{-1mm}\min_{\bm{\theta}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\mathcal{D}_{f}(\min\{\mathbb{P}+\delta,1\}||\max\{\mathbb{Q}-\delta,0\}),$ (12)
which is a regular minimization problem and (sub)gradient descent can be utilized to solve it.
@@ 4 Experiments
We use three popular notions of group fairness: demographic parity, equalized odds, and equality of opportunity violations (see Appendix A for definitions) to measure the fairness of trained models. To run Algorithm 1, we set $\eta_{{\bm{\theta}}}$ and $\eta_{\alpha}$ to $10^{-5}$ and $10^{-6}$ respectively in all experiments. Further, by changing $\lambda$, we get different points in the trade-off curve between accuracy and fairness. The range of $\lambda$ depends on the $f$-divergence (see Appendix J for more information on tuning hyper-parameters). In the inference phase of our experiments, we use the standard maximum likelihood decoding based on the output of the softmax layer, i.e., the predicted label is the label with the highest logit value.
@@ 4.1 Fairness-Accuracy Tradeoffs on Benchmark Datasets
Figure 1: Performance of different $f$-divergences as the regularizers. The experiment is on the adult dataset with gender and race as sensitive attributes. While the offered tradeoffs are close to each other for small demographic parity violations, KL-divergence shows an extraordinary performance for a low-fairness high-accuracy regime. We do not display the performance for larger batch sizes or when only one sensitive attribute is available due to the insignificant difference between the performance of different $f$-divergences.
In the first set of experiments, we compare different $f$-divergence formulations for ($f$-FERM) to each other and several state-of-the-art approaches supporting multiple sensitive attributes. Figure 1 demonstrates the given tradeoff on the adult dataset (Becker & Kohavi, 1996) with gender and race as the sensitive attributes (black-female, black-male, white-female, white-male). To measure fairness, we use the demographic parity violation defined as:
$\vspace{-1mm}\textrm{DPV}=\max_{i,j\in\mathcal{S}}|\mathbb{P}(\hat{y}=1|s=i)-\mathbb{P}(\hat{y}=1|s=j)|$
In the case of binary sensitive attributes (e.g., gender), there is no significant variation between different $f$-divergences. However, when we have $2$ sensitive attributes and the batch size is small ($8$ in Figure 1), the results significantly differ for various $f$-divergences. Interestingly, KL-divergence for smaller $\lambda$ values shows improvement in fairness violation and accuracy simultaneously. We do not observe such a phenomenon for other $f$-divergences and state-of-the-art approaches in the literature. Further, in Figure 2, we compare one of the $f$-divergences (reverse KL) to several SOTA methods including Mary et al. (2019); Baharlouei et al. (2020); Cho et al. (2020). Other approaches such as the pre-processing method of Zemel et al. (2013), post-processing approach of Hardt et al. (2016), and several in-processing methods including Zafar et al. (2017); Donini et al. (2018); Jiang et al. (2020) demonstrate lower performance compared to the ones depicted in Figure 2 and are removed from the figure. While our approach demonstrates consistently good performance across different batch sizes (full-batch, $64$, $8$, $2$), the performances of other methods drop significantly for smaller ones. For further experiments on other datasets (German and COMPAS) and other fairness measures (equality of opportunity and equalized odds violations), see Appendix K.
@@ 4.2 Fairness-Accuracy Tradeoffs in the Presence of the Distribution Shift
We perform two experiments to evaluate the Algorithms developed in Section 3. In the first experiment, we randomly switch the label of genders for $n\%$ of the data points ($n$ ranges from $1$ to $20$) in the Adult dataset. Then, we train models on the new datasets with a proportion of corrupted sensitive attributes and evaluate the performance on the test data. Figure 3 is obtained by training different models to achieve $80\%$ accuracy on the test data and comparing their demographic parity violation. By increasing the percentage of corrupted sensitive attributes, we see that both $f$-DRO and $f$-infinity achieve less DP violation than SOTA approaches in the literature. In this specific experiment, $f$-DRO works better than $f$-infinity, and there is no significant difference between choosing KL-divergence or $\chi^{2}$ as the function $f$. Among the papers designed for handling distribution shifts, Rezaei et al. (2021) and Wang et al. (2020) were the only options with the available implementation.
@@
In a more recently collected dataset (new adult) (Ding et al., 2021), the users are separated based on their living state. We train different fair models in a single state and evaluate the fairness-accuracy tradeoff in other states. Figure 4 depicts the performance of different methods. For each method, the center point is the average of accuracy and fairness among $50$ states. The horizontal and vertical lines show the $25$-percentile to $75$-percentile range of performance among the states. The training fairness violation is set to $0.02$ for all methods. We observe that $f$-infinity preserves the fairness level better than other approaches. In comparison, $f$-DRO has a better accuracy. Depending on the application, we suggest using $f$-infinity if preserving a high level of fairness is a priority and $f$-DRO for the cases when a better tradeoff between fairness and accuracy is expected. Note that both approaches offer better fairness-accuracy tradeoffs compared to the SOTA approaches in the literature.
@@
5 Conclusion
This paper presented a unified stochastic framework for fair empirical risk minimization via $f$-divergences ($f$-FERM). The key idea is to reformulate the objective function as a min-max optimization problem using Legendre-Fenchel duality of $f$-divergence. This enables us to develop an unbiased gradient estimator and a convergent stochastic first-order algorithm. Furthermore, we robustified $f$-FERM using $\ell_{p}$ norm balls as the uncertainty set against distributional changes. While our empirical investigation delves into the performance and fairness distinctions among various $f$-divergences, a more comprehensive analysis is warranted to determine the optimal $f$-divergence concerning the tradeoff between performance and fairness, faster convergence, and asymptotic behaviors. Furthermore, the distributionally robust formulation of fair empirical risk minimization and the advantages of each formulation can be explored beyond $f$-divergences as the measure of fairness violation and $\ell_{p}$ norm balls as uncertainty sets.
Acknowledgements
This work was supported by the NSF CAREER Award CCF2144985, the AFOSR Young Investigator Program Award FA9550-22-1-0192, a gift from the USC-Meta Center for Research and Education in AI, and a gift from Google.
@@ Appendix A $f$-FERM for other notions of group fairness
This section shows how we can use alternative notions of fairness, such as equality of opportunity of equalized odds (Hardt et al., 2016) instead of demographic parity violation in $f$-FERM.
@@
Therefore, we must add a regularizer per each class label to satisfy the equalized odds notion. Other notions of fairness can be used in this framework as long as they can be represented as the conditional independence between sensitive attributes, predictions, and labels (Castelnovo et al., 2022).
@@ Appendix B $f$-divergences for continuous sensitive attributes and target variables
In Section 2, we developed a framework for promoting fairness for classification problems where both target labels and sensitive attributes are discrete variables. Hence, we could efficiently solve the variational formulation that arose through the designing of unbiased estimators. However, it is not uncommon to find applications of f-divergence regularizers in practice that require either the sensitive features or the output variable to be continuous or both to be continuous parameters. In such cases, the summation over the respective variable is replaced by an integral over the probability distribution. The challenging aspect is calculating the variational form’s integral and trailing supremum in the continuous domain.
@@ Appendix C $f$-divergences cover well-known notions of fairness violation
Proposition C.1.
(Polyanskiy & Wu, 2022, Theorem 2.3) Let $f$ be a convex function from $\mathbb{R}^{+}$ to $\mathbb{R}$, such that $f$ is convex, $f(1)=0$, $f$ is strictly convex in a neighborhood of $1$. Then $\mathcal{D}_{f}(\mathbb{P}||\mathbb{Q})=0$, if and only if $P=Q$.
@@
Proposition C.2.
Exponential Rényi Mutual Information (ERMI) (Mary et al., 2019; Lowy et al., 2022) is an $f$-divergence with $f(t)=(t-1)^{2}$
Proof.
Exponential Rényi Mutual Information is defined as (Lowy et al., 2022):
@@
Proof.
Based on Propositions C.2, ERMI is an $f$-divergence with $f(t)=(t-1)^{2}$. Therefore, the proposition is an immediate result of Lemma 3 in (Lowy et al., 2022). ∎
Proposition C.4.
Rényi correlation (Baharlouei et al., 2020) can be upper bounded by the $f$-divergences for the choice of $f(t)=(t-1)^{2}$.
Proof.
Based on Propositions C.2, ERMI is an $f$-divergence with $f(t)=(t-1)^{2}$. Therefore, the proposition is an immediate result of Lemma 2 in (Lowy et al., 2022). ∎
Remark C.5.
Mutual Information as the measure of fairness violation (Cho et al., 2020) is a special case of $f$-divergences for the choice of KL-divergence $f(t)=t\log(t)$ in ($f$-FERM).
Appendix D Proof of Proposition 2.1
Lemma D.1.
@@
Proof.
Let $g$ be the convex conjugate of the function $f$ defined as:
@@
Since $f$ is a lower semi-continuous convex function, by Fenchel-Moreau theorem (Ioffe & Tihomirov, 2009), it is biconjugate, which means the taking conjugate of $g$ transforms it back to $f$. Therefore,
@@
where $g$ is the convex conjugate of $f$. ∎
@@
Set $\mathbb{P}=\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s}),\mathbb{Q}=\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s),$ and $\alpha_{i}=A_{jk}$. Therefore, we obtain the formulation in (3).
@@ Appendix E Derivation of Closed-Form Expressions for Unbiased Gradient Estimators of $f$-Divergences
Proof.
Proof follows naturally from (Polyanskiy & Wu, 2022, Proposition 7.2)
Theorem E.2.
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}A_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)-(A_{jk}+\frac{A_{jk}^{2}}{4})\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (19)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}A_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)+(1+\ln(-A_{jk}))\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (20)
@@
Theorem E.4.
Let $f(t)=\frac{1}{2}|t-1|$ and $\mathbb{P}(s=k)=\pi_{k}$ (Total Variational Distance). Then, Equation (1) can be written as (where $|A_{jk}|\leq\frac{1}{2}$):
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}A_{jk}\Big{[}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)-\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (21)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}A_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)-e^{A_{jk}-1}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (22)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}A_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)+\ln(2-e^{A_{jk}})\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{]}$ (23)
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}A_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)\\ -\frac{\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)}{\alpha}\Big{(}\Big{(}(\alpha-1)A_{jk}+1\Big{)}^{\frac{\alpha}{\alpha-1}}-1\Big{)}\Big{]}$ (24)
@@
This is equivalent to the substitution $\alpha\leftarrow\frac{1+\alpha}{2}$ in the original definition of generalized f-divergence.
Theorem E.8.
@@
$\min_{{\bm{\theta}}}\max_{{\mathbf{A}}}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j}\sum_{k}\pi_{k}\Big{[}A_{jk}\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j|s=k)+\mathbb{P}(\hat{y}_{{\bm{\theta}}}=j)\Big{(}\frac{1}{A_{jk}}+2\Big{)}\Big{]}$ (25)
@@
Appendix F Formal Statement of Theorem 2.2 and Proof
Theorem F.1.
@@
Now, we are ready to prove Theorem F.1.
Proof.
The proof consists of three main steps. First, we need to show that the gradient estimator in Algorithm 1 is unbiased. Since the samples are IID, for any function $\psi(\cdot,\cdot)$, and an IID batch of data points $\mathcal{B}$ we have:
@@
$\min_{{\bm{\theta}}}\>\max_{{\mathbf{A}}}\quad\frac{1}{n}\sum_{i=1}^{n}\left[\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}A_{jk}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}(A_{jk})\pi_{k}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\Big{]}\right]$ (27)